OpenStudy (anonymous):

Express in simplest radical form. √5/13

OpenStudy (anonymous):

\[\sqrt{5 \over 13}\]

OpenStudy (anonymous):

I don't think it can get any simpler

OpenStudy (lgbasallote):

\[\frac{\sqrt5}{\sqrt{13}}\] rationalize by multiplying numerator AND denominator by \(\sqrt{13}\)

sam (.sam.):

\[\sqrt{5 \over 13}=\frac{\sqrt5}{\sqrt{13}}=\frac{\sqrt5\sqrt{13}}{\sqrt{13}\sqrt{13}}=\frac{\sqrt{65}}{13}\]